Write in slope intercept form m=-1/5,(-2,1)

Found 2 solutions by Alan3354, helpplease@:
Answer by Alan3354(69443)

Makes no sense.

Answer by helpplease@(1)

y-y2=m (x-x2)
y-1=-1/5 (x+2)
y= -1/5x+0.6
Answer^